Determine the volume generated bounded by the curves y= 1/x, 2x - y =0, x=10, and the x-axis. Rotated about the x-axis. 


1
Expert's answer
2021-10-28T07:17:50-0400

According to second theorem of Guldino, volume obtined by a rotation of a section bounded by a function f(x) 

and the x-axis between a and b is :"V=\pi\int_a^bf^2(x)dx"


In our case, we have 

"V= \pi \int_1^2e^{(2x)}dx=\pi/2(e^4-e^1)"

so it will be: 


"2x-y=0\space"

"\u2014> 2(10)-y=0"

"Y=20m"